### v1

Abstract
We introduce HeadZoom, a hands-free interaction technique for navigating two-dimensional visual content using head movements. The system enables fluid zooming and panning by only using real-time head tracking. It supports natural control in applications such as map exploration, radiograph inspection, and image browsing, particularly where physical interaction is limited. We evaluated HeadZoom in a within-subjects user study comparing three interaction techniques—Static, Tilt Zoom, and Parallel Zoom—across spatial, error, and subjective metrics. Results show that Parallel Zoom significantly reduced total head movement compared to Static and Tilt modes. Users reported significantly lower perceived exertion for Parallel Zoom, confirming its suitability for prolonged or precision-based tasks. By minimizing movement demands while maintaining task effectiveness, HeadZoom advances the design of head-based 2D interaction in VR and creates new opportunities for immersive, accessible, and hands-free systems for image exploration.
keywords:
Head-Based Interaction, 2D Image Navigation, Hands-Free Interfaces, Zoom and Pan, MR User Interfaces.1 Introduction
Mixed Reality (MR) is moving beyond entertainment and into real-world domains, such as surgery, fieldwork, manufacturing, and education, where users’ hands are often occupied, gloved, sterilized, or constrained. In these contexts, reaching for a controller or touching a surface is inconvenient, impractical, unsafe, or impossible.
As MR technology advances, the demand for innovative hands-free interaction methods continues to grow, particularly in applications involving exploring and manipulating two-dimensional (2D) visual content such as maps, radiographs, and photographs. Traditional input devices, such as hand-held controllers or touchscreens, can be cumbersome or impractical in scenarios requiring hands-free interaction. To address these limitations, head-based interaction offers a promising alternative, enabling users to zoom in and pan across images using natural head movements.
HeadZoom introduces a novel framework for head-controlled zooming and panning of 2D content. The system uses precise head motion tracking to provide smooth and responsive control. This makes it particularly suitable for applications where precision and efficiency are critical, such as navigating geographical maps, examining medical radiographs, or exploring personal or professional photographs. This paper presents the development and evaluation of HeadZoom, with the following key contributions:
Head-Based Interaction Framework: A new method uses head movement to control the zoom and panning of 2D content, ensuring smooth transitions and accurate operation.
User-Centered Design: An interface that reduces cognitive and physical effort, with efficient interaction in extended use.
Extensive Evaluation: Comparing HeadZoom with existing interaction paradigms, demonstrating its effectiveness in exploring maps, analyzing radiographs, and browsing photographs.
By redefining how head motion is used for interaction, HeadZoom provides a powerful alternative for interacting with two-dimensional visual content, supporting advances in areas such as geographic information systems, medical imaging, and digital photography.
In the remainder of this paper, we first review related work and clarify the research gap that HeadZoom addresses and the novel contribution of the work. Next, we describe the implementation of HeadZoom, before presenting a user study to evaluate the technique. Section 7 contains the user study results discussed in Section 8. Finally, we close with conclusions and future research directions.
2 Related Work
Navigating visual content often requires panning (moving the viewport) and zooming (adjusting scale). Zooming and positioning techniques are critical to improving navigation efficiency in different applications. Many modern software tools, such as image editors or maps on the Web, provide interfaces for zooming- and panning-based navigation of visual content. This section reviews existing research in these areas and highlights the role of head-based interaction in addressing current challenges.
2.1 Zooming and Panning Interfaces
Zooming and panning interfaces have been actively investigated in 2D graphical user interfaces (GUI). Perlin and Fox [41, 4] pioneered using zooming and panning as a key method to navigate and manage an infinite 2D information space at various scale levels. They also introduced a concept of semantic zooming that presented a single entity in different representations depending on the zoom level.
Zooming and panning interfaces have been proposed for the 2D desktop GUI. You et al. [57] conducted a usability evaluation of the zoom and pan functions in web-based map interfaces, comparing different layouts of the zoom and pan buttons. Gutwin and Fedak [17] variants of panning and zooming interfaces to fit large 2D GUIs to smaller screens. Igarashi and Hinckley [23] proposed automatically controlling the zoom level proportional to the scrolling speed to help navigate large documents in a 2D GUI. A usability study showed that users can navigate through a large document efficiently and smoothly without getting disoriented, compared to manual control, which suffers from fast visual flow when scrolling too quickly while zoomed in. Cockburn et al. [13] provide a comprehensive review of early research on interfaces for navigating large documents in 2D desktop GUI, categorizing them into Overview+Detail, Zooming, Focus+Context, and Cue-based techniques. Standard input methods for zooming and panning in 2D GUI include mouse, touch gestures, and shortcut keys [57]. In addition to these popular techniques, researchers also investigated using other input methods, such as 6DOF controllers [45] or gaze input [51, 19].
The popularization of mobile devices with small screens further emphasized the need to explore zooming and panning techniques [17]. Although the pinch-to-zoom method became widely adopted with the arrival of modern touch screen-based mobile devices [1], its clear limitation of occupying both hands motivated many researchers to focus on providing a one-handed solution to control zooming and panning on mobile devices [33]. To overcome such limitations, researchers have developed various touch gestures [14, 27] to replace or complement the standard pinch-to-zoom gesture. At the same time, others also explored using additional detectors, such as pressure sensors combined with a touch screen to enable natural zooming and scrolling [35, 53], or using a gyro sensor or buttons built into the device [16].
Zooming and panning interfaces have also been explored in AR and VR applications. Typically, handheld controllers [11] or hand gestures and gaze input [47, 42, 34] are used to control zoom and pan in VR / AR applications, while some adopt touch-screen-based interaction [37, 10]. Lee et al. [28] proposed an automatic control of the zoom level based on the distance between the user’s view and the target object, increasing the zoom level as the AR camera approaches the tracking target, enabling users to have a closer look of the content with less tracking failure. While they linked the zoom control to the movement of the user’s view, the users held the AR camera in their hands. In comparison, our work investigates how head movement, as one of the hands-free input methods, could be useful for controlling zooming and panning operations in exploring large high-resolution images in VR.
2.2 Hands-Free Input Methods
Advancements in human-computer interaction have led to the development of alternative input methods beyond traditional keyboard and mouse interfaces. Hand gesture tracking as input modality has been actively developed and used as a natural interaction method for 3D interaction in both VR [46, 56] and AR [47, 6] environments. Despite its versatility and wide adoption in VR/AR interfaces, one of the typical limitations of hand gesture-based input is gorilla arm syndrome, where prolonged arm use leads to fatigue [40]. As an alternative, hands-free input methods [44, 30] have gained attention due to their potential to improve accessibility and usability. Hands-free interaction includes input modalities that do not require physical contact with a controller. Common approaches include voice commands, eye tracking, and head tracking [30].
Speech Control. Speech interfaces allow users to issue commands verbally or dictate text input [12, 31]. Speech interaction has been considered a natural and efficient interaction modality in XR systems [36, 43], especially in combination with gesture-based interaction [7, 29]. Although effective for discrete commands, speech is less suitable for continuous or highly granular adjustments [21, 24].
Eye Tracking. Eye tracking technology has been actively researched in the context of human-computer interaction [32, 49]. Gaze direction is a natural way to control an interface, as the eyes instinctively focus on areas of interest, and has also been actively investigated in VR [44, 42] and AR [26, 39] interfaces. Modern eye trackers using infrared corneal reflections achieve high accuracy [32, 49]. A key challenge, however, is the Midas touch problem, which makes it difficult to distinguish intentional interactions from normal eye movements [25, 32]. Although some modern VR/AR HMDs integrate eye tracking technology, the technology is mostly supported on high-end devices. It has not reached the status of being widely adopted as a standard feature.
Head-Based Interaction. Head tracking has been studied to control computers using natural head movements [22, 52]. Early work in human-computer interaction identified head movement as a simple and effective pointing method, often more accurate than voice or brain-computer interfaces [49]. It is also considered to be more affordable compared to eye tracking technology.
Head-based input is especially useful in assistive technology, allowing users with limited limb mobility to move a cursor or make selections by moving their head, using a webcam or infrared sensor. Systems such as Camera Mouse [5, 38] and Enable Viacam (eViacam)111https://eviacam.crea-si.com track the face of a user through a webcam, translating head movements into pointer control [8]. These systems have been used successfully by individuals with high spinal cord injuries, as head and eye movements remain intact [5, 49].
Head tracking is also a key enabler technology for inducing a sense of presence in VR [48, 55]. Modern VR/AR HMDs typically integrate head tracking technology to use head orientation and position to control the user’s viewpoint. Harrison and Dey’s Lean and Zoom system [20] detected users leaning toward or away from the screen, automatically magnifying the content. Users reported that this head-driven zooming felt natural and fluid, making the text easier to read and reducing discomfort.
While prior work has thoroughly explored zooming and panning techniques across desktop, mobile, and AR/VR platforms, most approaches rely on hand-held controllers, touch interfaces, or indirect mappings such as pressure sensors or gesture recognition. These methods often impose physical constraints or require significant setup, making them less suitable for hands-free continuous interaction in immersive contexts. Existing hands-free alternatives—such as gaze-based or voice-controlled interfaces—struggle with precision or lack suitability for granular control. Although head tracking has been used for cursor control and viewpoint manipulation, its potential for fine-grained, dual-purpose navigation (zoom-and-pan) in immersive 2D content exploration remains underexplored. Moreover, previous systems typically rely on singular head-motion mappings (e.g., lean-to-zoom) and lack comparative evaluations of multiple head-based techniques within the same task framework. To address this gap, we introduce and evaluate HeadZoom, systematically studying three distinct interaction techniques within a controlled experimental setting.
3 The HeadZoom Prototype
HeadZoom is a hands-free interaction framework that provides precise, controller-free zooming and panning in VR. As shown in Figure 2, the system continuously captures the HMD position and orientation, smoothing through Kalman filters to suppress jitter while preserving responsiveness. A virtual image plane is placed in front of the user, and raycasting from the filtered head pose determines the target location for interaction. This plane-based mapping enables natural, spatially consistent simultaneous panning and zooming using only head motion. The HMD’s position vector provides real-time 3D coordinates of the user’s head in space. This vector is used to compute the user’s distance from the image plane, which modulates the zoom level—closer positions result in a zoom-in effect, while retreating causes zoom-out. The forward vector represents the HMD directional orientation, specifically the normalized vector pointing where the user is facing. This vector is crucial for raycasting: It projects a ray from the user’s current position toward the image plane. The intersection of this ray with the plane determines the panning target, allowing the system to infer where in the image the person is attending to. Both vectors are filtered to reduce noise before being passed to the technique controller.
As illustrated in Figure 1, HeadZoom supports three head-driven interaction techniques that vary in how motion is mapped to image behavior. In the Static condition, zooming is achieved by moving closer to or farther from a fixed image in space, similar to real-world inspection. Panning is not actively supported; the image remains stationary regardless of head rotation.
In the Parallel condition, head translations toward the image plane modulate zoom level. At the same time, panning is enabled by shifting the image laterally in response to head rotation—mimicking motion across a virtual, front-facing plane. The image remains fixed in orientation, enhancing predictability during navigation.
The Tilt condition builds on Parallel by coupling the orientation of the image plane to the user’s head pose. The zooming behavior remains identical, but the panning dynamically reorients the image to maintain perpendicular alignment with the user’s view. This technique compensates for angular displacement and maintains a consistent visual relationship between user and content, potentially reducing rotational distortion during exploration.
4 Application Scenarios
Although the primary evaluation of HeadZoom focused on a controlled image search task, the system was designed with broader use cases in mind. Head-based interaction is particularly valuable in contexts where traditional input methods are impractical or unavailable. To demonstrate the system’s generalizability, we explored how HeadZoom could be applied across several domains requiring precise, hands-free navigation of two-dimensional content.
4.1 Applications in Specific Domains
To ensure the relevance and applicability of our proposed interaction techniques, we considered potential real-world applications where efficient and natural zooming mechanisms are critical. Our selection of application areas was guided by the need for high navigation precision, frequent zoom interactions, and varying user constraints, such as the availability of hands-free control. These application areas include the following:
Geographic Information Systems (GIS): GIS interfaces rely on frequent panning and zooming. Traditional GIS software employs mouse-based interaction, while mobile applications use multitouch gestures [57]. Researchers have also explored voice- and gesture-based interactions [15].
Medical Imaging: Radiologists and pathologists navigate large and detailed images, and traditional input devices such as mice can be inefficient. Alternatives such as head tracking, gesture control, and voice commands have been evaluated for image navigation [8, 54].
Digital Photography: Photo browsing is typically based on scrolling grids and pinch-to-zoom gestures. Some researchers investigated alternative approaches such as PhotoMesa, a zoomable image browser [3]. Head-directed control offers a hands-free alternative.
To explore the broader applicability of the HeadZoom system, we extended the experimental setup to include three domain-specific use cases: GIS, medical imaging, and digital photography. These scenarios demonstrate the system’s versatility across diverse tasks and user groups, moving beyond our controlled search experiment.
Geographic Information Systems (Google Maps): The application integrates dynamic map data rendered as 256×256 pixel tiles at various zoom levels. Users are tasked with locating iconic landmarks in globally recognized cities using only head-based input: leaning forward and backward adjusts the zoom level, while head rotations pan the map horizontally and vertically (see Figure 3.a).
However, the implementation revealed technical challenges. The core issue involved mismatched numerical precision: map coordinates are expressed in double precision, while headset tracking operates with single-precision floats. Even minimal head movements could result in abrupt positional jumps at higher zoom levels. Rapid gestures often led to unintended multi-step changes, causing the system to load new tile sets and introducing visible latency.
We implemented a ’focus zoom’ technique loosely inspired on Baudisch et al. [2]. During zoom-out motions, the map area within a circular region around the user’s gaze retains the current zoom level, while the surrounding map zooms out. The inverse is applied during zoom-in actions. This approach delays full tile updates until the user’s head stabilizes, reducing perceptual discontinuities. For panning, future refinements include lower-res zoom tiles as visual placeholders, while higher-resolution tiles load incrementally.
Medical Imaging (Chest X-Rays): HeadZoom can be adapted for radiological image inspection, using high-resolution chest X-rays obtained, for example, as 12-bit DICOM files from the MIMIC data set; cf. Figure 3.b. This use case targets professional radiologists and shifts the task paradigm from target localization to clinical evaluation, similarly to [50]. Navigation across image sets is achieved using a wireless button, while a Bluetooth mouse allows users to adjust windowing parameters critical for diagnostic visibility.
Crucially, this prototype incorporates eye tracking, allowing for the collection of gaze data throughout the task. A dedicated button toggles a heatmap overlay that visualizes gaze concentration over the image, supporting both post hoc analysis and real-time feedback.
Digital Photography (Image Browsing) The third application revisits image browsing in a consumer context. Inspired by earlier systems such as PhotoMesa [3], the application enables users to explore large photo collections using head-based zoom and pan controls. Unlike GIS and medical imaging setups, which prioritize analytical precision, this scenario emphasizes intuitive navigation and user experience. The interface presents image grids where users can seamlessly zoom in individual photos and pan across the collection cf. Figure 3.c, offering a hands-free alternative to traditional gesture-based browsing.
These three scenarios exemplify the adaptability of head-based interaction across domains with varied requirements, ranging from real-time responsiveness and visual continuity in GIS, to clinical accuracy and data logging in medical imaging, and user-centric exploration in photography.
5 Experiment
We conducted a controlled user study to evaluate the usability and effectiveness of our Head Zoom interaction technique in a digital photography scenario. The study aimed to assess participants’ performance and subjective experience in different zooming methods in Virtual Reality (VR) environments. We collected both quantitative measures (task completion time, success rate, head movement) and qualitative feedback to understand how each technique affects user engagement, task efficiency, and overall comfort.
5.1 Task
To reflect the panning and zooming demands in digital photography scenarios, we adopted Where’s Wally [18] task, which was also used in Brown et al. [9]’s 2D visual search study. Where’s Wally is a well-known visual search activity that is easy to explain but challenging to execute, as its large images require substantial panning and zooming. Moreover, this task provides images with varying difficulty levels, enabling us to select functions of different complexities to comprehensively evaluate our Head Zoom interaction technique.
We selected seven images from Where’s Wally. An image, featuring the fewest characters and minimal distractions, was used for training so that the participants could familiarize themselves with the three interaction scenarios (Static, Parallel, and Tilt). The remaining six images were classified into three low- and three medium-difficulty categories according to difficulty based on approximate character count and background complexity.
5.2 Task Image Characteristics and Visual Challenge
The pictures selected for the experiment came from the ”Where’s Wally?”222Where’s Wally? Images © Martin Handford / Walker Books Ltd. Used under fair use for academic research purposes. All images used in the study, including training and trial stimuli, are available in the supplementary materials accompanying this article. series and varied considerably in visual complexity.
Each image depicts a complex, cartoon-style environment populated by dozens to hundreds of small, visually similar characters, distributed across spatially and semantically varied backgrounds. Despite their playful appearance, these scenes pose significant cognitive and perceptual challenges for visual search tasks.
The training image, city, is the least complex of the set, featuring a modest number of characters with relatively clear segmentation and minimal foreground clutter. In contrast, the remaining six images vary widely in layout and visual load. Others, e.g. beach, introduce dense and unstructured clustering of figures, many partially occluded by umbrellas and beach props, which creates substantial camouflage effects. Other images are highly populated but exhibit more regular spatial organization.
The most visually demanding scenes combine extreme crowd density with intricate environmental elements, such as market stalls, fortress walls, snowy slopes, intense color variation, and overlapping figures. These images often present large contiguous areas of visual noise, making it difficult to distinguish targets based solely on spatial cues. The combination of character repetition, distractor similarity, and environmental detail results in high visual entropy, amplifying the difficulty of head-driven zoom and pan interaction.
This deliberate diversity in complexity ensured that the interaction techniques were evaluated across a realistic spectrum of search conditions, ranging from moderately challenging to perceptually saturated scenes.
5.3 Prototype Design
We developed our task prototype in Unity3D (version 6) for the Meta Quest 3 VR HMD. As shown in HeadZoom: Hands-Free Zooming and Panning for 2D Image Navigation Using Head Motion.c, each trial presented participants with a large Where’s Wally image (2800×1749 pixels). A red ring (approximately 105 pixels in radius) overlaid the image to indicate the current raycast hit point. In the upper left corner of the image, the participants saw the assigned interaction scenario (e.g. static, parallel, or tilt) and the trial number, ensuring that they were aware of the condition they were experiencing.
Before the first trial, participants calibrated the headset to establish how far forward and backward they could tilt their heads for maximum zoom‐in and zoom‐out. Their neutral head position was recorded as the central reference point. During the trial, participants used the designated interaction technique, moving their heads following these calibrated lean thresholds to zoom in or out. Panning occurred naturally through head rotation or subtle adjustments in body posture. Once Wally was centred in the red ring, participants pressed the controller’s trigger to indicate they had located the target.
Each trial allocated participants two minutes to locate Wally and allowed up to three attempts. A trial ended if participants successfully aligned Wally within the red ring, failed three times, or exceeded the two-minute limit.
All trial-related data, including video recordings, sensor readings, and image metrics, are logged in tab-separated CSV files. We also captured participants’ head orientation (pitch, yaw, roll) throughout each attempt for further analysis.
5.4 Experimental Design
Our study adopted a within subjects design for the participants that included three interaction scenarios (Static, parallel, and tilt) and two difficulty levels (low and medium). Figure 4.a shows an overview of the experimental timeline. Each participant was randomly assigned one low and one medium difficulty image under each scenario, resulting in six total trials (3 scenarios x 2 difficulty levels). We counterbalanced and randomized the order of both scenarios and difficulty levels to mitigate the potential bias from time‐sequence effects or subjective judgments.
Each scenario consisted of two trials (one low-difficulty image and one medium-difficulty image), presented in a counterbalanced sequence. Upon finishing each trial, participants completed a subjective rating (on a 1–7 scale) to indicate their perceived challenge in locating Wally within that image.
After completing both trials in a scenario, participants removed the headset and filled a post-scenario questionnaire assessing system usability and user experience. Upon concluding all three scenarios (i.e., six trials), they completed a final post-experience questionnaire, ranking each interaction technique in terms of zooming efficiency and object-finding accuracy.
5.5 Procedure
Figure 4.b illustrates the overall experimental setup. Upon arrival, participants were given an information sheet describing the procedures, followed by completion of a consent form and a pre‐experience questionnaire. Then, we configured the Quest headset and seated participants in a fixed‐position chair (Figure 4.c). We used Meta Horizon to stream the headset’s view for monitoring participant performance. Participants then underwent a 15 minute training session, familiarising themselves with all three interaction scenarios for the Where’s Wally task. Participants then proceeded according to the process depicted in 5.4. We used Google Forms to collect all pre-, post‐scenario, and post‐experience responses. The experiment lasted approximately 40 minutes, after which each participant received a $20 gift card as compensation.
5.6 Participants
We recruited 31 participants (16 women, 15 men, ages 19 to 49). A total of 8 participants (10. 5%) reported never using VR, 19 participants (61.3%) had fewer than 5 hours of VR experience, 1 (3.2%) reported having 5–20 hours of experience, 8 participants (25.8%) had logged more than 20 hours, and 3 participants (9.7%) had never used VR. Among all participants, 8 (25.8%) indicated they had never experienced VR‐related motion sickness, 4 (12.9%) were unsure or had insufficient VR experience, and 19 (61.3%) experienced it only occasionally.
6 Metrics
To assess user performance and understand the impact of each interaction technique, we computed quantitative metrics from the headset data logs, grouped into four categories: Time Completion, Spatial Metrics, Error‐Related Metrics, and Subjective Difficulty.
6.1 Time Completion Metrics
This category evaluates how quickly and successfully participants locate Wally. The main metric is Completion Time, which measures the time taken to correctly identify Wally.
6.2 Spatial Metrics
These metrics quantify physical interactions and exploration patterns, including head movement, rotation, lean, zoom usage, and participant preferences.
Total Head Movement. Is is the cumulative headset displacement where represents headset position at frame .:
Total Head Rotation. Summation of angular displacement between consecutive forward vectors :
Lean Usage. Measures maximum lean from the initial headset position across all frames.
Hover Time on Characters. Time spent, gazing near a specific character (Wally, Wenda, Wizard, Odlaw):
with binary indicator:
Zoom Usage. Evaluates how participants interact with the zoom functionality. These metrics help characterize participants’ zoom interaction behaviors, reflecting how actively they used zoom capabilities. They are measured through different metrics:
Zoom Change Count counts how many times participants meaningfully change the zoom level during the trial.
Total Zoom Distance corresponds to the cumulative amount participants zoomed in and out across the entire trial.
Average Zoom indicates participants’ average zoom level throughout the trial. Maximum Zoom records the highest level of magnification achieved during the trial.
6.3 Error‐Related Metrics
These metrics reflect errors made during task execution:
False Positives. Count of incorrect identification attempts.
Task Success Rate. A binary metric indicating task success (finding Wally under 120 seconds and within three tries).
6.4 Subjective Difficulty Metrics
Participants rated each image’s difficulty using a 1–7 scale (ImageUserGrade), where 1 = ”very difficult” and 7 = ”very easy.” This subjective assessment allows correlation analysis with objective performance metrics (e.g., Completion Time, False Positives), highlighting discrepancies or consistencies between perceived difficulty and actual performance.
7 Results
To evaluate the performance and interaction characteristics of the three interaction techniques (Static, Parallel, and Tilt), we performed a comprehensive statistical analysis on data collected from a controlled VR experiment involving 31 participants. The task involved participants in locating a target character (Wally) within images of varying complexity. The statistical evaluation process consisted of three main stages and is presented in Figure 5:
Repeated Measures Analysis of Variance (ANOVA) was used to identify general differences between the three interaction techniques for each metric. We used repeated measures ANOVA rather than one-way ANOVA because each participant experienced all three interaction conditions (Static, Parallel, and Tilt), creating within-subject dependencies in the data. This approach accounts for individual participant differences and provides greater statistical power by reducing error variance. We considered the differences statistically significant at a value .
Post-hoc pairwise comparisons were made using paired t-tests to pinpoint specific differences between interaction techniques for metrics identified as significant by ANOVA. To address the issue of multiple comparisons that inflate Type I errors, we applied the Bonferroni correction, adjusting the significance threshold to ( divided by comparisons per metric).
Effect sizes were calculated using Cohen’s d to quantify the significant differences. Cohen’s d was calculated and used to compute Effect sizes. We interpret these as negligible , small , medium , or large .
7.1 Interaction Pattern Differences
The ANOVA results revealed significant differences among interaction modes for out of metrics analyzed, with F-statistics ranging from to (all ). The most pronounced differences were observed for Average Head Rotation (, ) and Average Head Movement (, ).
After the Bonferroni correction, out of pairwise comparisons remained statistically significant. The key differences were:
Static Mode required significantly more head movement (, ) and rotation (, ) compared to Parallel mode (both ). Static mode also showed significantly higher average lean () and maximum zoom (), indicating greater user physical effort.
Parallel Mode exhibited the least head movement () and rotation (), with significantly fewer zoom changes (), indicating a more constrained and efficient interaction pattern.
Tilt Mode generally showed intermediate values for most physical metrics, indicating moderate physical interaction compared to the other two modes.
Figure 7 shows the normalized averaged performance metrics for Total Head Movement, Max Zoom, Total Zoom Distance, Zoom Change Count and Max Lean, revealing that Parallel is the most ergonomic mode, since it leads to lower physical movements. Figure 8 presents the distributions of the Total Head Rotation per interaction mode, indicating that the Static mode required more movements when compared to the other interaction modes.
Effect sizes further clarify the magnitude of these differences, highlighting the large effect sizes between Static and Parallel for AvgHeadRotation () and AvgHeadMovement (). Similarly, we observed large effect sizes between Parallel and Tilt for AvgHeadRotation () and between Static and Tilt modes for AvgHeadRotation () and AvgHeadMovement ().
7.2 Task Performance and Efficiency
Despite significant differences in physical interaction patterns, no statistically significant differences emerged regarding task completion time (, ). Mean completion times were similar across modes: Static with secs (), Parallel with seconds (), and Tilt with seconds (). Additionally, no significant differences were found for false positive rates or subjective difficulty ratings. This suggests that all three interaction techniques provided similar efficiency and accuracy, despite differences in physical interaction demands. Figure 6 shows the distribution of the completion times across participants.
7.3 Usability Evaluation
The System Usability Scale (SUS) scores, assessing subjective usability, did not show significant differences between interaction modes (). All three modes achieved high usability scores: Static (), Parallel (), and Tilt ().
These scores indicate excellent usability for all interaction methods. The analysis of individual SUS items also revealed that there were no significant differences in perceived complexity, ease of use, learnability, confidence, or support needs, confirming that differences in physical interaction did not significantly influence overall usability perceptions.
7.4 User Preferences
The Parallel mode was most frequently selected as the preferred method, with the highest ”best” ratings, while Static and Tilt modes were more frequently identified as the least preferred. Specifically, Parallel was chosen most frequently as the ”best” interaction mode by participants and less often marked as the ”worst,” contrasting with Tilt, which received the highest number of ”worst” ratings. This subjective preference aligns with objective findings on reduced physical demand for the Parallel mode, underscoring user appreciation for its ergonomic efficiency.
7.5 Impact of Image Complexity
“Time On” metrics (TimeOnWally, TimeOnWenda, TimeOnWizard, TimeOnOdlaw) indicate the duration participants focused on specific targets or distractors did not differ significantly across interaction modes. This finding held across varying image complexities, suggesting consistent interaction patterns regardless of visual clutter.
7.6 Gender Differences
In terms of gender, only the User Difficulty Rating showed significant gender differences, with males rating tasks as slightly easier () compared to women (). Further analysis revealed a significant gender-interaction mode effect for SUS scores specifically in the Parallel mode, where males reported significantly higher usability scores () compared to females (), indicating gender may influence subjective experience particularly in this interaction condition.
8 Discussion
The following section discusses the implications and significance of our main experimental findings. Each finding is contextualized within the broader scope of VR interaction research, highlighting its relevance and potential impact on VR interface design and usability.
Finding 1: Significant Differences in Physical Interaction Patterns. Our study revealed significant differences in physical interaction patterns among the three modes (Static, Parallel, and Tilt), particularly regarding head movement, rotation, and zoom usage. The Static mode had significantly higher physical engagement, evidenced by greater head movement and rotation, than the Parallel mode. Conversely, the Parallel mode demonstrated minimal physical exertion requirements, making it the most ergonomically efficient. This aligns with the participants’ subjective preferences, as the Parallel mode was frequently the preferred interaction mode and less frequently the least preferred. The Tilt mode generally showed moderate values, indicating that this technique balances physical engagement and comfort. This finding is important, as it suggests that interaction techniques significantly affect user comfort and could influence user adoption, especially in prolonged VR applications or precision-demanding tasks like medical imaging analysis.
Finding 2: Comparable Task Performance Across Interaction Modes. Despite pronounced physical differences among interaction techniques, no significant differences were observed in task completion times, accuracy, or subjective difficulty ratings. This indicates that while the techniques differ significantly in ergonomic efficiency, they do not affect the effectiveness or accuracy of task execution. These results suggest that designers of VR interfaces have flexibility in choosing interaction techniques based on considerations other than raw performance metrics, such as user comfort, accessibility, or specific application requirements. Thus, the decision regarding interaction technique selection can prioritize user comfort and reduce physical strain without compromising task performance.
Finding 3: High Usability Scores Independent of the Interaction Technique. All interaction modes received high usability ratings, with no significant differences observed in the overall SUS scores or in individual usability dimensions such as complexity, ease of use, learnability and confidence. This suggests that despite differences in how users physically interacted with the VR system, all interaction methods provided an excellent user experience. High usability scores across all modes imply that users readily adapted to different interaction paradigms, reinforcing that usability perceptions are robust to variations in physical interaction patterns. Consequently, one can confidently employ varied interaction methods tailored to specific contextual or ergonomic considerations without concern for negative usability impacts.
Finding 4: Gender Differences in Interaction Experiences. Our analysis identified notable gender-related differences, specifically in subjective task difficulty ratings and usability perceptions within the Parallel interaction mode. Male participants rated the tasks as slightly easier and reported significantly higher usability for the Parallel mode than female participants. These gender differences highlight the importance of considering demographic factors in interaction design, suggesting that user groups might perceive certain interaction methods differently.
Finding 5: Robustness of Observed Differences. Applying the Bonferroni correction minimally impacted the significant differences observed in our analysis, with only a minor reduction from 16 to 15 significant pairwise comparisons. This highlights the robustness of our findings, indicating that the reported differences in physical interaction patterns and ergonomic efficiency are reliable and unlikely to be Type I errors. The large effect sizes observed, particularly between Static and Parallel modes, further substantiate the practical significance of these differences. These robust results reinforce the recommendation of the Parallel mode as an ergonomic choice for extended or precision-based VR tasks.
Our findings underscore the importance of ergonomic considerations in VR interaction design, demonstrate flexibility in achieving comparable performance outcomes across interaction methods, and highlight user preferences and demographic differences as critical considerations for user-centered VR system development.
9 Design Guidelines
Based on the user study results and our findings, there are some guidelines that could be provided for people developing hands-free, controller-free VR interaction, including:
G1: Minimize effort: User interface techniques should be developed that favor mappings with low head movement requirements (e.g., Parallel) to reduce fatigue.
G2: Stabilize input with filtering: Use smoothing filters to suppress jitter in continuous control tasks.
G3: Ensure spatial consistency: Maintain predictable image behavior to reduce disorientation (as seen in Tilt).
G4: Support Visual Anchoring: Design interactions that allow users to maintain visual focus during input. Avoid mappings that disrupt line-of-sight or require large shifts in viewpoint.
G5: Preserve visual focus: Avoid techniques that disrupt gaze alignment during navigation.
G6: Balance control and simplicity: Choose interaction modes that generalize well and require minimal learning.
10 Limitations and Future Research
HeadZoom advances the field of spatial interaction by enabling fluid, continuous panning and zooming through natural head movement. Although promising, this approach also introduces new challenges and opportunities. HeadZoom demonstrates the feasibility of hands-free head-based interaction for 2D content navigation, but several limitations remain. First, the system relies solely on head pose as input, which constrains the granularity of control and may not scale well to tasks that require high-precision manipulation or sustained focus over extended sessions. In the future it would be good to explore other input alternatives, especially multimodal options, such as combining eye-gaze and head poose.
Another limitation is that the currently implemented system assumes stable seating posture and uniform viewing distance, which may not hold in ambulatory or multi-user VR contexts. In the future we should conduct experiments supporting a wider range of physical movement.
Finally, the research was conducted with relatively simple tasks. In order for the results to generalize, future studies should be conducted with a wider range of tasks. This will help establish how applicable technique is in a wider range of VR settings.
Future Research should address the following directions:
Ergonomics and Fatigue: Prolonged head movement may cause neck strain or discomfort for the user. Investigating ergonomic thresholds and incorporating adaptive mechanisms could mitigate fatigue and improve long-term usability [40].
Precision and Responsiveness: Head-based input must support fine-grained control and minimal latency for effective navigation. Further studies should examine how system responsiveness affects task accuracy and user satisfaction [15].
Multimodal Integration: Combining head-based control with complementary modalities, e.g. eye gaze or speech, may improve precision, reduce workload, and support diverse user needs [49].
Application-Specific Optimization: Future implementations should tailor head-based navigation to domain-specific requirements, including image granularity in medical imaging, map scale in GIS, and browsing patterns in photography [8].
By addressing these directions, future systems can harness the full potential of head-based interaction while ensuring usability, comfort, and domain adaptability.
11 Conclusions
This paper presented HeadZoom, a hands-free interaction technique that enables users to pan and zoom two-dimensional imagery using natural head motion. Through a controlled user study involving three interaction conditions: static, parallel, and tilt, we demonstrated that head-based control can be effective and usable in diverse image navigation tasks.
Among the tested methods, the Parallel mode consistently minimized physical effort while maintaining task performance and was rated most favorably by participants. Statistical analysis confirmed significant reductions in head movement and zoom-related actions, suggesting its suitability for applications requiring sustained or precise visual exploration. Although the Static condition yielded slightly faster average times, it imposed greater physical demand, particularly in rotational metrics. Although promising in theory, the Tilt condition was often perceived as disorienting due to its dynamic image realignment.
Our findings underscore the feasibility of replacing traditional input modalities with controller-free embodied techniques in mixed reality contexts. Future research should explore multi-modal extensions (e.g., gaze, speech) and domain-specific adaptations for clinical, cartographic, and creative workflows. In addition, more research is needed on adaptive calibration strategies and long-term ergonomics to enhance robustness and inclusivity among broader user populations.
Supplemental Materials
Supplemental materials include:
The questionnaires filled by Subjects during the experiments
a copy of all the images used during the user tests. The dataset includes seven images. One was used for training (resized_city) and the remaining six were used for training.
a companion video
These are available as additional items in the precisionconference submission
Acknowledgements.
The authors wish to thank A, B, and C. This work was supported in part by a grant from XYZ.References
- [1] J. Avery, M. Choi, D. Vogel, and E. Lank. Pinch-to-zoom-plus: an enhanced pinch-to-zoom that reduces clutching and panning. In Proceedings of the 27th annual ACM symposium on User interface software and technology, pp. 595–604, 2014.
- [2] P. Baudisch, N. Good, and P. Stewart. Focus plus context screens: combining display technology with visualization techniques. In Proceedings of the 14th Annual ACM Symposium on User Interface Software and Technology, UIST ’01, p. 31–40. ACM, 2001. doi: 10.1145/502348.502354
- [3] B. Bederson and J. Hollan. Photomesa: A zoomable image browser. ACM Transactions on Graphics, 23(3):501–508, 2004. doi: 10.1145/1015706.1015756
- [4] B. B. Bederson, J. D. Hollan, K. Perlin, J. Meyer, D. Bacon, and G. Furnas. Pad++: A zoomable graphical sketchpad for exploring alternate interface physics. Journal of Visual Languages & Computing, 7(1):3–32, 1996.
- [5] M. Betke, J. Gips, and P. Fleming. The camera mouse: visual tracking of body features to provide computer access for people with severe disabilities. IEEE Transactions on neural systems and Rehabilitation Engineering, 10(1):1–10, 2002.
- [6] M. Billinghurst, T. Piumsomboon, and H. Bai. Hands in space: Gesture interaction with augmented-reality interfaces. IEEE computer graphics and applications, 34(1):77–80, 2014.
- [7] R. A. Bolt. “put-that-there” voice and gesture at the graphics interface. In Proceedings of the 7th annual conference on Computer graphics and interactive techniques, pp. 262–270, 1980.
- [8] A. Brown and R. White. Head-tracking as an interface device for image control in digital pathology: a comparative study. Medical Image Analysis, 38:115–130, 2021. doi: 10.1016/j.media.2021.102104
- [9] E. T. Brown, A. Ottley, H. Zhao, Q. Lin, R. Souvenir, A. Endert, and R. Chang. Finding waldo: Learning about users from their interactions. IEEE Transactions on Visualization and Computer Graphics, 20(12):1663–1672, 2014. doi: 10.1109/TVCG.2014.2346575
- [10] W. Büschel, A. Mitschick, T. Meyer, and R. Dachselt. Investigating smartphone-based pan and zoom in 3d data spaces in augmented reality. In Proceedings of the 21st International Conference on Human-Computer Interaction with Mobile Devices and Services, pp. 1–13, 2019.
- [11] H. Chang and M. F. Cohen. Panning and zooming high-resolution panoramas in virtual reality devices. In Proceedings of the 30th Annual ACM Symposium on User Interface Software and Technology, pp. 279–288, 2017.
- [12] L. Clark, P. Doyle, D. Garaialde, E. Gilmartin, S. Schlögl, J. Edlund, M. Aylett, J. Cabral, C. Munteanu, J. Edwards, et al. The state of speech in hci: Trends, themes and challenges. Interacting with computers, 31(4):349–371, 2019.
- [13] A. Cockburn and A. Karlson. A review of zoomable user interfaces: Foundations and future directions. ACM Computing Surveys, 40(2):5:1–5:39, 2008. doi: 10.1145/1357054.1357213
- [14] M. Farhad and I. S. MacKenzie. Evaluating tap-and-drag: A single-handed zooming method. In International Conference on Human-Computer Interaction, pp. 233–246. Springer, 2018.
- [15] D. Foster and J. Lopez. Gesture-based map interface: Exploring hands-free navigation. Geospatial Computing, 5(1):98–112, 2021. doi: 10.1109/GEO.2021.4521032
- [16] S. Garg and I. Scott MacKenzie. Comparison of touch and touchless zoom control methods for single-handed mobile interaction. In Advances in Human Factors and System Interactions: Proceedings of the AHFE 2021 Virtual Conference on Human Factors and Systems Interaction, July 25-29, 2021, USA, pp. 71–78. Springer, 2021.
- [17] C. Gutwin and C. Fedak. Interacting with big interfaces on small screens: a comparison of fisheye, zoom, and panning techniques. In Proceedings of Graphics Interface 2004, pp. 145–152, 2004.
- [18] M. Handford. Where’s Wally? Walker Books, London, 1987.
- [19] D. W. Hansen, H. H. Skovsgaard, J. P. Hansen, and E. Møllenbach. Noise tolerant selection by gaze-controlled pan and zoom in 3d. In Proceedings of the 2008 symposium on Eye tracking research & applications, pp. 205–212, 2008.
- [20] C. Harrison and A. Dey. Lean and zoom: Proximity-aware user interface for touchscreen devices. In Proceedings of CHI, pp. 1003–1012, 2008. doi: 10.1145/1357054.1357213
- [21] D. Hepperle, Y. Weiß, A. Siess, and M. Wölfel. 2d, 3d or speech? a case study on which user interface is preferable for what kind of object interaction in immersive virtual reality. Computers & Graphics, 82:321–331, 2019.
- [22] M. Hunke and A. Waibel. Face locating and tracking for human-computer interaction. In Proceedings of 1994 28th Asilomar Conference on Signals, Systems and Computers, vol. 2, pp. 1277–1281. IEEE, 1994.
- [23] T. Igarashi and K. Hinckley. Speed-dependent automatic zooming for browsing large documents. In Proceedings of the 13th annual ACM symposium on User interface software and technology, pp. 139–148, 2000.
- [24] T. Igarashi and J. F. Hughes. Voice as sound: using non-verbal voice input for interactive control. In Proceedings of the 14th Annual ACM Symposium on User Interface Software and Technology, UIST ’01, p. 155–156. ACM, 2001. doi: 10.1145/502348.502372
- [25] R. J. Jacob. What you look at is what you get: eye movement-based interaction techniques. In Proceedings of the SIGCHI conference on Human factors in computing systems, pp. 11–18, 1990.
- [26] M. Kytö, B. Ens, T. Piumsomboon, G. A. Lee, and M. Billinghurst. Pinpointing: Precise head-and eye-based target selection for augmented reality. In Proceedings of the 2018 CHI conference on human factors in computing systems, pp. 1–14, 2018.
- [27] J. Lai, D. Zhang, and S. Wang. Contextzoom: a single-handed partial zooming technique for touch-screen mobile devices. International Journal of Human–Computer Interaction, 33(6):475–485, 2017.
- [28] G. A. Lee, H. Bai, and M. Billinghurst. Automatic zooming interface for tangible augmented reality applications. In Proceedings of the 11th ACM SIGGRAPH International Conference on Virtual-Reality Continuum and its Applications in Industry, pp. 9–12, 2012.
- [29] M. Lee, M. Billinghurst, W. Baek, R. Green, and W. Woo. A usability study of multimodal input in an augmented reality environment. Virtual Reality, 17:293–305, 2013.
- [30] M. Lee and E. Zhang. Development of a sensing platform based on hands-free interfaces for controlling electronic devices. Sensors, 22(6):3098, 2022. doi: 10.3390/s22063098
- [31] R. Lu, R. Wei, and J. Zhang. Human-computer interaction based on speech recognition. Applied and Computational Engineering, 36:102–110, 2024.
- [32] P. Majaranta and A. Bulling. Eye tracking and eye-based human–computer interaction. In Advances in physiological computing, pp. 39–65. Springer, 2014.
- [33] D. Mendes, M. Sousa, A. Ferreira, and J. Jorge. Thumbcam: Returning to single touch interactions to explore 3d virtual environments. In Proceedings of the Ninth ACM International Conference on Interactive Tabletops and Surfaces, ITS ’14, p. 403–408. ACM, 2014. doi: 10.1145/2669485.2669554
- [34] A. Messaci, N. Zenati, M. Belhocine, and S. Otmane. Zoom-fwd: Efficient technique for 3d gestual interaction with distant and occluded objects in virtual reality. Computer Animation and Virtual Worlds, 33(1):e2033, 2022.
- [35] T. Miyaki and J. Rekimoto. Graspzoom: zooming and scrolling control model for single-handed mobile interaction. In Proceedings of the 11th International Conference on Human-Computer Interaction with Mobile Devices and Services, pp. 1–4, 2009.
- [36] J. Muller, C. Krapichler, L. S. Nguyen, K. H. Englmeier, and M. Lang. Speech interaction in virtual reality. In Proc. of the 1998 IEEE Intl. Conf. on Acoustics, Speech and Signal Processing, ICASSP’98 (Cat. No. 98CH36181), vol. 6, pp. 3757–3760. IEEE, 1998.
- [37] A. Mulloni, A. Dünser, and D. Schmalstieg. Zooming interfaces for augmented reality browsers. In Proceedings of the 12th international conference on Human computer interaction with mobile devices and services, pp. 161–170, 2010.
- [38] M. Nabati and A. Behrad. Camera mouse implementation using 3d head pose estimation by monocular video camera and 2d to 3d point and line correspondences. In 2010 5th International Symposium on Telecommunications, pp. 825–830. IEEE, 2010.
- [39] H. M. Park, S. H. Lee, and J. S. Choi. Wearable augmented reality system using gaze interaction. In 2008 7th IEEE/ACM International Symposium on Mixed and Augmented Reality, pp. 175–176. IEEE, 2008.
- [40] S. Patterson and O. Chen. Study on gorilla arm fatigue in mid-air gestural interfaces. Purdue University News, 2017. doi: 10.1007/s10209-017-00712-x
- [41] K. Perlin and D. Fox. Pad: an alternative approach to the computer interface. In Proceedings of the 20th annual conference on Computer graphics and interactive techniques, pp. 57–64, 1993.
- [42] K. Pfeuffer, B. Mayer, D. Mardanbegi, and H. Gellersen. Gaze+pinch interaction in virtual reality. In Proceedings of the 5th symposium on spatial user interaction, pp. 99–108, 2017.
- [43] T. Piumsomboon, D. Altimira, H. Kim, A. Clark, G. Lee, and M. Billinghurst. Grasp-shell vs gesture-speech: A comparison of direct and indirect natural interaction techniques in augmented reality. In 2014 IEEE International Symposium on Mixed and Augmented Reality (ISMAR), pp. 73–82. IEEE, 2014.
- [44] T. Piumsomboon, G. Lee, R. W. Lindeman, and M. Billinghurst. Exploring natural eye-gaze-based interaction for immersive virtual reality. In 2017 IEEE symposium on 3D user interfaces (3DUI), pp. 36–39. IEEE, 2017.
- [45] T. Robinson and L. Green. A comparative study of input devices for digital slide navigation. Journal of Digital Imaging, 33:1234–1245, 2020. doi: 10.1007/s10278-020-00311-4
- [46] K. M. Sagayam and D. J. Hemanth. Hand posture and gesture recognition techniques for virtual reality applications: a survey. Virtual Reality, 21:91–107, 2017.
- [47] K. A. Satriadi, B. Ens, M. Cordeil, B. Jenny, T. Czauderna, and W. Willett. Augmented reality map navigation with freehand gestures. In 2019 IEEE conference on virtual reality and 3D user interfaces (VR), pp. 593–603. IEEE, 2019.
- [48] M. Slater. Immersion and the illusion of presence in virtual reality. British journal of psychology, 109(3):431–433, 2018.
- [49] J. Smith and J. Doe. Eye tracking and head movement detection: A state-of-art survey. Journal of Human-Computer Interaction, 35(4):567–590, 2020. doi: 10.1007/s10209-020-00718-w
- [50] M. Sousa, D. Mendes, S. Paulo, N. Matela, J. Jorge, and D. S. o. Lopes. Vrrrroom: Virtual reality for radiologists in the reading room. In Proceedings of the 2017 CHI Conference on Human Factors in Computing Systems, CHI ’17, p. 4057–4062. ACM, 2017. doi: 10.1145/3025453.3025566
- [51] S. Stellmach and R. Dachselt. Investigating gaze-supported multimodal pan and zoom. In Proceedings of the Symposium on Eye Tracking Research and Applications, pp. 357–360, 2012.
- [52] P. Strumiłło and T. Pajor. A vision-based head movement tracking system for human-computer interfacing. In 2012 Joint Conference New Trends In Audio & Video and Signal Processing: Algorithms, Architectures, Arrangements and Applications (NTAV/SPA), pp. 143–147. IEEE, 2012.
- [53] K. Suzuki, R. Sakamoto, D. Sakamoto, and T. Ono. Pressure-sensitive zooming-out interfaces for one-handed mobile interaction. In Proceedings of the 20th International Conference on Human-Computer Interaction with Mobile Devices and Services, pp. 1–8, 2018.
- [54] H. Wilson and L. Cooper. A gesture-based tool for sterile browsing of radiology images. IEEE Transactions on Medical Imaging, 38(8):1975–1984, 2019. doi: 10.1109/TMI.2019.2907764
- [55] T. L. Wu, A. Gomes, K. Fernandes, and D. Wang. The effect of head tracking on the degree of presence in virtual reality. International Journal of Human–Computer Interaction, 35(17):1569–1577, 2019.
- [56] L. Yang, J. Huang, T. Feng, W. Hong-An, and D. Guo-Zhong. Gesture interaction in virtual reality. Virtual Reality & Intelligent Hardware, 1(1):84–112, 2019.
- [57] M. You, C.-w. Chen, H. Liu, and H. Lin. A usability evaluation of web map zoom and pan functions. International Journal of Design, 1(1), 2007.